What type of solution do you get for quadratic equations where D < 0? Give reasons for your answer. Also provide an example of such a quadratic equation and find the solution of the equation.

Answer by stanbon(75887)

What type of solution do you get for quadratic equations where D < 0? Give reasons for your answer. Also provide an example of such a quadratic equation and find the solution of the equation.
----------
If the discriminant is negative you get two distinct complex solutions.
---------------------
y = x^2+x+1
x = [-1 +- sqrt(1-4*1*1)]/2
x = (-1 + sqrt(-3))/2 or x = (-1-sqrt(-3))/2
x = (-1+isqrt(3))/2 or x = (-1-isqrt(3))/2
====================
Cheers,
Stan H.